# Regularized Online RLHF with Generalized Bilinear Preferences

### v1

@@ Abstract
We consider the problem of contextual online RLHF with general preferences, where the goal is to identify the Nash Equilibrium. We adopt the Generalized Bilinear Preference Model (GBPM) to capture potentially intransitive preferences via low-rank, skew-symmetric matrices. We investigate general preference learning with any strongly convex regularizer (where $\eta^{-1}$ is the regularization strength), generalizing beyond prior works limited to reverse KL-regularization. Central to our analysis is proving that the dual gap of the greedy policy is bounded by the square of the estimation error—a result derived solely from strong convexity and the skew-symmetricity of GBPM. Building on this insight and a feature diversity assumption, we establish two regret bounds via two simple algorithms: (1) Greedy Sampling achieves polylogarithmic, $e^{{\mathcal{O}}(\eta)}$-free regret $\tilde{{\mathcal{O}}}(\eta d^{4}(\log T)^{2})$. (2) Explore-Then-Commit achieves $\mathrm{poly}(d)$-free regret $\tilde{{\mathcal{O}}}(\sqrt{\eta rT})$ by exploiting the low-rank structure; this is the first statistically efficient guarantee for online RLHF in high-dimensions.
@@
Table 1: Comparison of regret bounds for online RLHF under GBPM; for unregularized regret, take $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\eta$$\infty$. For simplicity, we set $\delta=\frac{1}{T}$ but note that all guarantees hold with high probability. Also, we consider only $\mathrm{poly}(d,C_{\min}^{-1})$-dependencies.
@@ 2.2 Generalized Bilinear Preference Model (GBPM)
Let rank $r\leq\lfloor d/2\rfloor$ and norm bound $S>0$ be known. For any ${\bm{x}}\in{\mathcal{X}},{\bm{a}}^{1},{\bm{a}}^{2}\in{\mathcal{A}}$, the ground-truth preference is: $P^{*}({\bm{a}}^{1}\succ{\bm{a}}^{2}\mid{\bm{x}}):=\mu\left(\phi({\bm{x}},{\bm{a}}^{1})^{\top}\bm{\Theta}_{\star}\phi({\bm{x}},{\bm{a}}^{2})\right),$ where $\bm{\Theta}_{\star}\in\mathrm{Skew}(d;2r,S)$ is an unknown skew-symmetric, low-rank matrix, and $\mu:\mathbb{R}\rightarrow[0,1]$ is a known link satisfying: for some $\kappa>0$, 1. $\mu$ is twice differentiable, monotone increasing, and symmetric ($\mu(z)+\mu(-z)=1$). 2. $\kappa\leq\dot{\mu}\left(\phi^{\top}\bm{\Theta}\phi^{\prime}\right)\leq L_{\mu},\ \left|\ddot{\mu}\left(\phi^{\top}\bm{\Theta}\phi^{\prime}\right)\right|\leq L_{\mu},\ \ \forall\bm{\phi},\bm{\phi}^{\prime}\in{\mathcal{B}}^{d}(1),\bm{\Theta}\in\mathrm{Skew}(d;2r,S)$. [^2] We assume the same constant $L_{\mu}$ for the upper bounds of $\dot{\mu}$ and $|\ddot{\mu}|$ for simplicity of the exposition.
@@ 2.3 Regularized NE and Regret Definitions
For any $\bm{\Theta}\in\mathrm{Skew}(d)$, the value of the max-min game, $\max_{\pi^{1}}\min_{\pi^{2}}J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(\pi^{1},\pi^{2};\bm{\Theta})$, is always $\frac{1}{2}$.
@@
For any $\{({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}^{1}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\hat{\pi}_{t}^{2}})\}_{t\in[T]}$, (a) Average-Best-Response Regret: $\mathrm{ABR\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T):=\sum_{t=1}^{T}\left\{\mathrm{DGap}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}^{1}})+\mathrm{DGap}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\hat{\pi}_{t}^{2}})\right\}.$ (b) Max-Best-Response Regret: $\mathrm{MBR\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T):=\sum_{t=1}^{T}\mathrm{DGap}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}^{1}}).$
@@ 3 A New Analysis of Regularized Regret
For any estimator $\widehat{\bm{\Theta}}_{t}\in\mathrm{Skew}(d)$ at time $t$, define the max-player’s policy as $\hat{\pi}_{t}\leftarrow\operatorname*{arg\,max}_{\pi^{1}}\min_{\pi^{2}}J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(\pi^{1},\pi^{2};\widehat{\bm{\Theta}}_{t}).$ (7) Then, the instantaneous dual gap is simultaneously bounded as follows: denoting ${\bm{E}}_{t}:=\bm{\Theta}_{\star}-\widehat{\bm{\Theta}}_{t}$, $\mathrm{DGap}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}})$ $\leq(L_{\mu}^{2}{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}\beta+L_{\mu})\mathbb{E}_{\bm{\phi}\sim{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}}}\left[\left\lVert{\bm{E}}_{t}\bm{\phi}\right\rVert_{2}^{2}\right],$ $\mathrm{DGap}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}})$ $\leq\frac{L_{\mu}}{2}\mathbb{E}_{\bm{\phi}\sim{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}}}\left[\left\lVert{\bm{E}}_{t}\bm{\phi}\right\rVert_{2}+\left\lVert{\bm{E}}_{t}\bm{\phi}\right\rVert_{2}^{2}\right].$
@@
For $\hat{\pi},\tilde{\pi}\in\Pi$ and $\bm{\Theta}_{\star},{\bm{E}}\in\mathrm{Skew}(d)$, $\left|\mathbb{E}\left[\dot{\mu}(\bm{\phi}^{\top}\bm{\Theta}_{\star}\tilde{\bm{\phi}})\bm{\phi}^{\top}{\bm{E}}\tilde{\bm{\phi}}\right]\right|\leq\frac{L_{\mu}}{2}{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\left\lVert\hat{\pi}-\tilde{\pi}\right\rVert_{1}}\mathbb{E}_{\bm{\phi}}[\left\lVert{\bm{E}}\bm{\phi}\right\rVert_{2}].$
@@
$J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(\hat{\pi},\pi)-J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(\hat{\pi},\tilde{\pi})\geq\langle\nabla_{\tilde{\pi}}J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(\hat{\pi},\tilde{\pi}),\pi-\tilde{\pi}\rangle+\frac{1}{2{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}\beta}{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\left\lVert\pi-\tilde{\pi}\right\rVert_{1}^{2}}.$
@@ 4.1 Norm + Skew-Constrained MLE
Let $\delta\in(0,1)$ and suppose that $d^{2}\log\frac{T}{d}\gtrsim\kappa^{-1}\log\frac{1}{\delta}$. Then, with probability at least $1-\delta,$ GS simultaneously attains the following bounds: $\mathrm{MBR\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T)$ $\lesssim{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}\beta\kappa^{-1}d^{4}C_{\min}^{-1}\left(\log\frac{T}{d}\right)^{2},$ $\mathrm{MBR\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T)$ $\lesssim\kappa^{-\frac{1}{2}}C_{\min}^{-\frac{1}{2}}d^{2}\sqrt{T}\log\frac{T}{d}.$
@@ 5.1 Nuclear-Norm Regularized MLE and Regret Bound
Let $\delta\in(0,1)$, and suppose $T\gtrsim\kappa_{\star}^{-2}C_{\min}^{-4}dr\log\frac{d}{\delta}$. Then, ETC with $\lambda_{T_{0}}=\sqrt{\frac{32L_{\mu}\log(4d/\delta)}{T_{0}}}$ attains $\mathrm{MBR\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T)\lesssim T_{0}$ with probability at least $1-\delta$, where $T_{0}$ is chosen as: $T_{0}\asymp\kappa^{-1}C_{\min}^{-2}\sqrt{T{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}\beta r\log\frac{d}{\delta}}$; or $T_{0}\asymp\left(\kappa^{-2}C_{\min}^{-4}rT^{2}\log\frac{d}{\delta}\right)^{1/3}$.
@@ C.1 Main Proof
Let $\delta\in(0,1)$, and $\tilde{\kappa}\in\{\kappa,\kappa_{\star}\}$ is such that $\tilde{\kappa}=\kappa_{\star}$ if 2 holds, and $\tilde{\kappa}=\kappa$ if otherwise. Suppose that $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}T_{0}\gtrsim\tilde{\kappa}^{-2}C_{\min}^{-4}dr\log\frac{d}{\delta}$}. Then, with $\lambda_{T_{0}}=\sqrt{\frac{32L_{\mu}\log\frac{4d}{\delta}}{T_{0}}}$, we have the following: ${\mathbb{P}}\left(\left\lVert\widehat{\bm{\Theta}}-\bm{\Theta}_{\star}\right\rVert_{F}\lesssim\frac{1}{\tilde{\kappa}C_{\min}^{2}}\sqrt{\frac{L_{\mu}r\log\frac{d}{\delta}}{T_{0}}}\right)\geq 1-\delta.$

### v2

# arXiv:2602.23116v2[cs.LG] 27 Feb 2026

@@ Abstract
We consider the problem of contextual online RLHF with general preferences, where the goal is to identify the Nash Equilibrium. We adopt the Generalized Bilinear Preference Model (GBPM) to capture potentially intransitive preferences via low-rank, skew-symmetric matrices. We investigate general preference learning with any strongly convex regularizer and regularization strength $\eta^{-1}$, generalizing beyond prior work limited to reverse KL-regularization. Central to our analysis is proving that the dual gap of the greedy policy is bounded by the square of the estimation error, a result derived solely from strong convexity and the skew-symmetry of GBPM. Building on this insight and a feature diversity assumption, we establish two regret bounds via two simple algorithms: (1) Greedy Sampling achieves polylogarithmic, $e^{{\mathcal{O}}(\eta)}$-free regret $\tilde{{\mathcal{O}}}(\eta d^{4}(\log T)^{2})$. (2) Explore-Then-Commit achieves $\mathrm{poly}(d)$-free regret $\tilde{{\mathcal{O}}}(\sqrt{\eta rT})$ by exploiting the low-rank structure; this is the first statistically efficient guarantee for online RLHF in high-dimensions.
@@
Table 1: Comparison of regret bounds for online RLHF under GBPM; for unregularized regret, take $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta$$\infty$. For simplicity, we set $\delta=\frac{1}{T}$ but note that all guarantees hold with high probability. Also, we consider only $\mathrm{poly}(d,C_{\min}^{-1})$-dependencies.
@@ 2.2 Generalized Bilinear Preference Model (GBPM)
Let rank $r\leq\lfloor d/2\rfloor$ and norm bound $S>0$ be known. For any ${\bm{x}}\in{\mathcal{X}},{\bm{a}}^{1},{\bm{a}}^{2}\in{\mathcal{A}}$, the ground-truth preference under GBPM is:
$P^{*}({\bm{a}}^{1}\succ{\bm{a}}^{2}\mid{\bm{x}}):=\mu\left(\phi({\bm{x}},{\bm{a}}^{1})^{\top}\bm{\Theta}_{\star}\phi({\bm{x}},{\bm{a}}^{2})\right),$
where $\bm{\Theta}_{\star}\in\mathrm{Skew}(d;2r,S)$ is an unknown skew-symmetric, low-rank matrix, and $\mu:\mathbb{R}\rightarrow[0,1]$ is a known link satisfying: for some $\kappa>0$,
  1. 1.
    $\mu$ is twice differentiable, monotone increasing, and symmetric ($\mu(z)+\mu(-z)=1$).
  2. 2.
    $\kappa\leq\dot{\mu}\left(\phi^{\top}\bm{\Theta}\phi^{\prime}\right)\leq L_{\mu},\ \left|\ddot{\mu}\left(\phi^{\top}\bm{\Theta}\phi^{\prime}\right)\right|\leq L_{\mu},\ \ \forall\bm{\phi},\bm{\phi}^{\prime}\in{\mathcal{B}}^{d}(1),\bm{\Theta}\in\mathrm{Skew}(d;2r,S)$. [^2] We assume the same constant $L_{\mu}$ for the upper bounds of $\dot{\mu}$ and $|\ddot{\mu}|$ for simplicity of the exposition.
@@ 2.3 Regularized NE and Regret Definitions
For any $\bm{\Theta}\in\mathrm{Skew}(d)$, the value of the game, $\max_{\pi^{1}}\min_{\pi^{2}}J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(\pi^{1},\pi^{2};\bm{\Theta})$, is always $\frac{1}{2}$.
@@
For any $\{({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}^{1}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\hat{\pi}_{t}^{2}})\}_{t\in[T]}$,
(a) Average-Best-Response Regret:
$\mathrm{ABR\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T):=\sum_{t=1}^{T}\left\{\mathrm{DGap}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}^{1}})+\mathrm{DGap}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\hat{\pi}_{t}^{2}})\right\}.$
(b) Max-Best-Response Regret:
$\mathrm{MBR\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T):=\sum_{t=1}^{T}\mathrm{DGap}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}^{1}}).$
@@ 3 A New Analysis of Regularized Regret
For any estimator $\widehat{\bm{\Theta}}_{t}\in\mathrm{Skew}(d)$ at time $t$, define the max-player’s policy as
$\hat{\pi}_{t}\leftarrow\operatorname*{arg\,max}_{\pi^{1}}\min_{\pi^{2}}J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(\pi^{1},\pi^{2};\widehat{\bm{\Theta}}_{t}).$ (7)
Then, the instantaneous dual gap is simultaneously bounded as follows: denoting ${\bm{E}}_{t}:=\bm{\Theta}_{\star}-\widehat{\bm{\Theta}}_{t}$,
$\mathrm{DGap}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}})$ $\leq(L_{\mu}^{2}{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}\beta+L_{\mu})\mathbb{E}_{\bm{\phi}\sim{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}}}\left[\left\lVert{\bm{E}}_{t}\bm{\phi}\right\rVert_{2}^{2}\right],$
$\mathrm{DGap}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}})$ $\leq\frac{L_{\mu}}{2}\mathbb{E}_{\bm{\phi}\sim{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\pi}_{t}}}\left[\left\lVert{\bm{E}}_{t}\bm{\phi}\right\rVert_{2}+\left\lVert{\bm{E}}_{t}\bm{\phi}\right\rVert_{2}^{2}\right].$
@@
For $\hat{\pi},\tilde{\pi}\in\Pi$ and $\bm{\Theta}_{\star},{\bm{E}}\in\mathrm{Skew}(d)$,
$\left|\mathbb{E}\left[\dot{\mu}(\bm{\phi}^{\top}\bm{\Theta}_{\star}\tilde{\bm{\phi}})\bm{\phi}^{\top}{\bm{E}}\tilde{\bm{\phi}}\right]\right|\leq\frac{L_{\mu}}{2}{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\left\lVert\hat{\pi}-\tilde{\pi}\right\rVert_{1}}\mathbb{E}_{\bm{\phi}}[\left\lVert{\bm{E}}\bm{\phi}\right\rVert_{2}].$
@@
[…]
$\geq\langle\nabla_{\tilde{\pi}}J_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(\hat{\pi},\tilde{\pi}),\pi-\tilde{\pi}\rangle+\frac{1}{2{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}\beta}{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\left\lVert\pi-\tilde{\pi}\right\rVert_{1}^{2}}.$
@@ 4.1 Norm + Skew-Constrained MLE
Let $\delta\in(0,1)$ and suppose that $d^{2}\log\frac{T}{d}\gtrsim\kappa^{-1}\log\frac{1}{\delta}$. Then, with probability at least $1-\delta,$ GS simultaneously attains the following bounds:
$\mathrm{MBR\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T)$ $\lesssim{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}\beta\kappa^{-1}d^{4}C_{\min}^{-1}\left(\log\frac{T}{d}\right)^{2},$
$\mathrm{MBR\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T)$ $\lesssim\kappa^{-\frac{1}{2}}C_{\min}^{-\frac{1}{2}}d^{2}\sqrt{T}\log\frac{T}{d}.$
@@ 5.1 Nuclear-Norm Regularized MLE and Regret Bound
Let $\delta\in(0,1)$, and suppose $T\gtrsim\kappa_{\star}^{-2}C_{\min}^{-4}dr\log\frac{d}{\delta}$. Then, ETC with $\lambda_{T_{0}}=\sqrt{\frac{32L_{\mu}\log(4d/\delta)}{T_{0}}}$ attains $\mathrm{MBR\text{-}Reg}_{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}(T)\lesssim T_{0}$ with probability at least $1-\delta$, where $T_{0}$ is chosen as:
  • $T_{0}\asymp\kappa^{-1}C_{\min}^{-2}\sqrt{T{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\eta}\beta r\log\frac{d}{\delta}}$; or
  • $T_{0}\asymp\left(\kappa^{-2}C_{\min}^{-4}rT^{2}\log\frac{d}{\delta}\right)^{1/3}$.
@@ B.1 Main Proof
Let $\tilde{\bm{\phi}}\sim\rho$ be such that $\mathbb{E}_{\tilde{\bm{\phi}}\sim\rho}\left[\tilde{\bm{\phi}}\tilde{\bm{\phi}}^{\top}\right]\succeq C_{\min}{\bm{I}}_{d}.$ Then, for any positive semi-definite ${\bm{M}}\in{\mathbb{R}}^{d^{2}\times d^{2}}$ and any vector $\bm{\phi}\in{\mathbb{R}}^{d}$, the following holds:
@@ C.1 Main Proof
Let $\delta\in(0,1)$, and $\tilde{\kappa}\in\{\kappa,\kappa_{\star}\}$ is such that $\tilde{\kappa}=\kappa_{\star}$ if 2 holds, and $\tilde{\kappa}=\kappa$ if otherwise. Suppose that $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}T_{0}\gtrsim\tilde{\kappa}^{-2}C_{\min}^{-4}dr\log\frac{d}{\delta}$}. Then, with $\lambda_{T_{0}}=\sqrt{\frac{32L_{\mu}\log\frac{4d}{\delta}}{T_{0}}}$, we have the following:
${\mathbb{P}}\left(\left\lVert\widehat{\bm{\Theta}}-\bm{\Theta}_{\star}\right\rVert_{F}\lesssim\frac{1}{\tilde{\kappa}C_{\min}^{2}}\sqrt{\frac{L_{\mu}r\log\frac{d}{\delta}}{T_{0}}}\right)\geq 1-\delta.$